# Splitting obstructions and $\mathbb{Z}_{2}$ invariants in time-reversal symmetric topological insulators

### v1

@@ 3. The topological invariant
Then we can use Lemma A.1 to fix a quaternionic basis of $\operatorname{Ran}(P(0,0))$, i.e., a basis compatible with the quaternionic structure induced by the time-reversal operator $T$: fix an orthonormal set $\{u_{j}\}_{j\in\{1,\cdots n\}}$, and define $u_{n+j}=Tu_{j}$ for $j\in\{1,\cdots,n\}$ to have a complete basis of $\operatorname{Ran}(P(0,0))$. Then we can get a continuous collection of orthonormal bases of $P(t,k_{2})$ by setting, for all $t\in[-\pi,\pi],k\in\mathbb{T}^{1},j\in\{1,\cdots,2n\}$
@@
where $J$ is as in (8). The aperiodicity in $k_{1}$ is expressed by a continuous family of matching matrices $\alpha(k_{2})\in U(2n)$ such that
$u_{j}(-\pi,k_{2})=\sum_{a=1}^{2n}[\alpha(k_{2})]_{j,a}u_{a}(\pi,k_{2}).$
This family satisfies the symmetry constraint $J=\overline{\alpha(k_{2})}J\alpha(-k_{2})$ because:
$\begin{split}\sum_{a=1}^{2n}[J]_{j,a}u_{a}(\pi,-k_{2})&=Tu_{j}(-\pi,k_{2})=T\sum_{a=1}^{2n}[\alpha(k_{2})]_{j,a}u(\pi,k_{2})\\ &=\sum_{a,b=1}^{2n}\overline{[\alpha(k_{2})]_{j,a}}[J]_{a,b}u_{b}(-\pi,-k_{2})=\\ &=\sum_{a,c=1}^{2n}\left[\overline{\alpha(k_{2})}(J)\right]_{j,a}[\alpha(-k_{2})]_{a,c}u_{c}(\pi,-k_{2})=\\ &=\sum_{a=1}^{2n}\left[\overline{\alpha(k_{2})}(J)\alpha(-k_{2})\right]_{j,a}u_{a}(\pi,-k_{2})\end{split}$
@@
Using this symmetry constraint we can find two matrices $\gamma_{0},\gamma_{\pi}$ such that $\alpha(0)=J\gamma_{0}^{t}J^{-1}\gamma_{0}$ and same for $\gamma_{\pi}$. For example, since $J\alpha(0)^{t}J^{-1}=\alpha(0)$, we can find a self-adjoint matrix $L$ such that $\alpha(0)=e^{iL}$ and $JL^{t}J^{-1}=L$, then $\gamma_{0}=e^{iL/2}$ will work. It is obvious that $\det(\alpha(0))=\det(\gamma_{0})^{2}$ and $\det(\alpha(\pi))=\det(\gamma_{\pi})^{2}$. If $\det(\gamma_{0})=e^{i\lambda_{0}}$ and $\det(\gamma_{\pi})=e^{i\lambda_{\pi}}$ for two real numbers $\lambda_{0},\lambda_{\pi}\in\mathbb{R}$, we can choose a continuous function $\mu:[0,\pi]\to\mathbb{R}$ such that $\det(\alpha(k_{2}))=e^{i\mu(k_{2})}$ and $\mu(0)=2\lambda_{0}$. Then it holds that $e^{i\mu(\pi)}=e^{i2\lambda_{\pi}}$, so $\mu(\pi)-2\lambda_{\pi}\in 2\pi\mathbb{Z}$.
@@
$\delta(P)=e^{i\frac{\left(2\lambda_{\pi}-\mu(\pi)\right)}{2}}\in\{\pm 1\}=\mathbb{Z}_{2}.$
@@
The quantity defined previously does not depend on the choices of $\lambda_{0},\lambda_{\pi}$ $\gamma_{0},\gamma_{\pi}$.
@@
Instead if we took another $\gamma^{\prime}_{0}$ such that $\alpha(0)=J\gamma_{0}^{t}J^{-1}\gamma_{0}=J(\gamma^{\prime}_{0})^{t}J^{-1}\gamma^{\prime}_{0}$, then it is true that:
$(\gamma^{\prime}_{0})^{t}J\gamma^{\prime}_{0}=\gamma_{0}^{t}J\gamma_{0}\Leftrightarrow J=(\gamma_{0}(\gamma_{0}^{\prime})^{-1})^{t}J(\gamma_{0}(\gamma_{0}^{\prime})^{-1})$
@@
This means that $\gamma_{0}(\gamma^{\prime}_{0})^{-1}\in\mathrm{Sp}(2n)$, so $\det(\gamma_{0}(\gamma^{\prime}_{0})^{-1})=1$ and therefore $\det(\gamma_{0})=\det(\gamma^{\prime}_{0})$. Since the same holds for $\gamma_{\pi}$, we have that $\det(\gamma_{0}),\det(\gamma_{\pi})$ depends only on $\alpha(0),\alpha(\pi)$. ∎
The invariant can also be formulated as the parity of a Chern number. In particular, if $P^{\pm}:\mathbb{T}^{2}\to\operatorname{Proj}_{n}(\mathcal{H})$ is a pair of projection-valued maps such that $TP^{+}(k)=P^{-}(-k)T$ and with $P(k)=P^{+}(k)+P^{-}(k)$, then $\delta(P)=e^{i\pi\operatorname{Ch}(P^{-})}$. However, this formulation requires an entire section of the paper to be fully explored. Nonetheless, we can borrow the main result of the next section to prove the following Proposition.

### v2

@@ 4.1. Review: time-reversal symmetric matching matrices
Then we can use Lemma A.1 to fix a quaternionic basis of $\operatorname{Ran}P(0,0)$, i.e., a basis compatible with the quaternionic structure induced by the time-reversal operator $T$: fix an orthonormal set $\{u_{j}\}_{j\in\{1,\cdots n\}}$, and define $u_{n+j}=Tu_{j}$ for $j\in\{1,\cdots,n\}$ to have a complete basis of $\operatorname{Ran}P(0,0)$. Then we can get a continuous collection of orthonormal bases of $P(t,k_{2})$ by setting, for all $t\in[-\pi,\pi]$, $k\in\mathbb{T}^{1}$, and $j\in\{1,\cdots,2n\}$,
@@
where $J$ is as in (12). The lack of periodicity in $k_{1}$ is expressed by a continuous family of matching matrices $\alpha(k_{2})\in\mathrm{U}(2n)$ such that
$u_{j}(\pi,k_{2})=\sum_{a=1}^{2n}[\alpha(k_{2})]_{j,a}u_{a}(-\pi,k_{2}).$
This family satisfies the symmetry constraint $J=\overline{\alpha(k_{2})}J\alpha(-k_{2})$:
$\begin{split}\sum_{a=1}^{2n}[J]_{j,a}u_{a}(-\pi,k_{2})&=Tu_{j}(\pi,-k_{2})=T\sum_{a=1}^{2n}[\alpha(k_{2})]_{j,a}u(-\pi,-k_{2})\\ &=\sum_{a,b=1}^{2n}\overline{[\alpha(k_{2})]_{j,a}}[J]_{a,b}u_{b}(\pi,k_{2})\\ &=\sum_{a,c=1}^{2n}\left[\overline{\alpha(k_{2})}(J)\right]_{j,a}[\alpha(-k_{2})]_{a,c}u_{c}(-\pi,k_{2})\\ &=\sum_{a=1}^{2n}\left[\overline{\alpha(k_{2})}(J)\alpha(-k_{2})\right]_{j,a}u_{a}(-\pi,k_{2})\end{split}$
Remark 4.1.
From the above symmetry constraint, we immediately deduce that, as stated in [panati2007triviality, monaco2015symmetry], any $P:\mathbb{T}^{2}\to\operatorname{Proj}_{2n}(\mathcal{H})$ is a time-reversal symmetric projection-valued map has $\operatorname{Ch}(P)=0$. Indeed, we can compute the Chern number through a symmetric family of matching matrices with $J=\overline{\alpha(k_{2})}J\alpha(-k_{2})$. This implies that $\det(\alpha(k_{2}))=\det(\alpha(-k_{2}))$, but a map with this property has a winding number equal to zero due to Lemma A.4.
One natural question arises: how will the matching matrix change after the choice of a different symmetric and a frame $\{v_{j}(t,k_{2})\}_{j\in\{1,\cdots,2n\}}$ which is periodic in the second entry but not in the first? Clearly, there is a unitary-valued map $\beta:[-\pi,\pi]\times\mathbb{T}^{1}\to\mathrm{U}(2n)$ such that $v_{j}(t,k_{2})=\sum_{a=1}^{2n}[\beta(t,k_{2})]_{j,a}u_{a}(t,k_{2})$. The first condition on $\beta$ is that the second frame is symmetric if and only if $J\beta(t,k_{2})=\overline{\beta(-t,-k_{2})}J$:
Moreover, if the new family of matching matrices is $\alpha^{\prime}(k_{2})$ with
$v_{j}(\pi,k_{2})=\sum_{a=1}^{2n}[\alpha^{\prime}(k_{2})]_{j,a}v_{a}(-\pi,k_{2}),$
then $\alpha$ and $\alpha^{\prime}$ are intertwined by the equation
(8) $\alpha^{\prime}(k_{2})=\beta(\pi,k_{2})\alpha(k_{2})\beta(-\pi,k_{2})^{-1}\quad\forall k_{2}\in\mathbb{T}^{1}.$
This is true because
$\begin{split}\sum_{a=1}^{2n}[\alpha^{\prime}(k_{2})]_{j,a}v_{a}(-\pi,k_{2})&=v_{j}(\pi,k_{2})\Leftrightarrow\\ \sum_{a=1}^{2n}[\alpha^{\prime}(k_{2})\beta(-\pi,k_{2})]_{j,a}u_{a}(-\pi,k_{2})&=\sum_{a=1}^{2n}[\beta(\pi,k_{2})]_{j,a}u_{a}(\pi,k_{2})\Leftrightarrow\\ \sum_{a=1}^{2n}[\alpha^{\prime}(k_{2})\beta(-\pi,k_{2})]_{j,a}u_{a}(-\pi,k_{2})&=\sum_{a=1}^{2n}[\beta(\pi,k_{2})\alpha(k_{2})]_{j,a}u_{a}(-\pi,k_{2}).\end{split}$
We can interpret the previous conditions as follows: the collection of pseudo-periodic symmetric frames for a time-reversal symmetric projection-valued map is in bijection with the set $A/\sim$, where
$A:=\left\{\alpha:\mathbb{T}^{1}\to\mathrm{U}(2n):J=\overline{\alpha(k_{2})}J\alpha(-k_{2})\right\}$
and where $\alpha\sim\alpha^{\prime}$ in $A$ if and only if there is a map $\beta:[-\pi,\pi]\times\mathbb{T}^{1}\to\mathrm{U}(2n)$ such that
(9) $J\beta(t,k_{2})=\overline{\beta(-t,-k_{2})}J\quad\mbox{and}\quad\alpha^{\prime}(k_{2})=\beta(\pi,k_{2})\alpha(k_{2})\beta(-\pi,k_{2})^{-1}\quad\forall k_{2}\in\mathbb{T}^{1},t\in[-\pi,\pi].$
Remark 4.2.
The first of the above two conditions implies that the determinant of $\beta(t,\cdot)$ has an even winding number for all $t\in[-\pi,\pi]$. In fact, the symmetry condition $J\beta(0,k_{2})=\overline{\beta(0,-k_{2})}J$, forces $\beta(0,0),\beta(0,\pi)\in\mathrm{Sp}(2n)$, which in turn implies that $\det\beta(0,0)=\det\beta(0,\pi)=1$. Moreover, $\det\beta(0,k_{2})=\overline{\det\beta(0,-k_{2})}$, so we can use Lemma A.3 to conclude that the winding number in the second direction $\det\beta(0,\cdot)]$ must be even. Continuity in $t$ ensures that $\det\beta(t,\cdot)]$ is even (and constant) for every $t\in[-\pi,\pi]$.
4.2. The $\mathbb{Z}_{2}$-valued invariant
The set of symmetric matching matrices has a topological invariant dating back to the milestone paper [kane2005z], where Kane and Mele demonstrated that it is possible to associate a $\mathbb{Z}_{2}$-valued topological invariant to a specific discrete honeycomb model. The methods introduced in that work were subsequently extended to a broader class of models in several later papers, including [fu2007topological, freed2013twisted, graf2013bulk, de2015classification, fiorenza2016z, Prodan_2009], each providing a distinct formulation of the invariant and a different interpretation of the physical information it encodes. In this subsection, elaborating on [Cornean_2017], we will present yet another formulation of the invariant, using the matching matrices formalism.
Consider a family of symmetric matching matrices $\alpha(k_{2})$ such that $J=\overline{\alpha(k_{2})}J\alpha(-k_{2})$ for all $k_{2}\in\mathbb{T}^{1}$. Using this symmetry constraint we can find two unitary matrices $\gamma_{0},\gamma_{\pi}$ such that
$\alpha(0)=J\gamma_{0}^{t}J^{-1}\gamma_{0}$
@@ 4.2. The $\mathbb{Z}_{2}$-valued invariant
and similarly for $\gamma_{\pi}$. For example, since $J\alpha(0)^{t}J^{-1}=\alpha(0)$, we can find a self-adjoint matrix $L$ such that $\alpha(0)=e^{iL}$ and $JL^{t}J^{-1}=L$: then $\gamma_{0}=e^{iL/2}$ will work. It is obvious that $\det\alpha(0)=(\det\gamma_{0})^{2}$ and $\det\alpha(\pi)=(\det\gamma_{\pi})^{2}$. If $\det\gamma_{0}=e^{i\lambda_{0}}$ and $\det\gamma_{\pi}=e^{i\lambda_{\pi}}$ for two real numbers $\lambda_{0},\lambda_{\pi}\in\mathbb{R}$, we can choose a continuous function $\mu:[0,\pi]\to\mathbb{R}$ such that $\det\alpha(k_{2})=e^{i\mu(k_{2})}$ and $\mu(0)=2\lambda_{0}$. Then it holds that $e^{i\mu(\pi)}=e^{i2\lambda_{\pi}}$, so $\mu(\pi)-2\lambda_{\pi}\in 2\pi\mathbb{Z}$.
@@
$\delta(P)=\delta(\alpha):=e^{i\frac{\left(2\lambda_{\pi}-\mu(\pi)\right)}{2}}\in\{\pm 1\}=\mathbb{Z}_{2}.$
@@
The quantity defined previously does not depend on the choices of $\gamma_{0},\gamma_{\pi}\in\mathrm{U}(2n)$, of $\lambda_{0},\lambda_{\pi}\in\mathbb{R}$, and is also well-posed with respect to the equivalence relation in (9), meaning that $\delta(\alpha)=\delta(\alpha^{\prime})$ if $\alpha\sim\alpha^{\prime}$ in $A$. Therefore it depends only the symmetric projection-valued map $P$, allowing for the abuse of notation $\delta(P)=\delta(\alpha)$.
@@
Instead if we took another $\gamma^{\prime}_{0}$ such that $\alpha(0)=J\gamma_{0}^{t}J^{-1}\gamma_{0}=J(\gamma^{\prime}_{0})^{t}J^{-1}\gamma^{\prime}_{0}$, then it is true that
$(\gamma^{\prime}_{0})^{t}J\gamma^{\prime}_{0}=\gamma_{0}^{t}J\gamma_{0}\Leftrightarrow J=(\gamma_{0}(\gamma_{0}^{\prime})^{-1})^{t}J(\gamma_{0}(\gamma_{0}^{\prime})^{-1}).$
@@
This means that $\gamma_{0}(\gamma^{\prime}_{0})^{-1}\in\mathrm{Sp}(2n)$, so $\det(\gamma_{0}(\gamma^{\prime}_{0})^{-1})=1$ and therefore $\det(\gamma_{0})=\det(\gamma^{\prime}_{0})$. Since the same holds for $\gamma_{\pi}$, we have that $\det(\gamma_{0}),\det(\gamma_{\pi})$ depends only on $\alpha(0),\alpha(\pi)$.
Instead, suppose that we want to study the invariant of the family
$\alpha^{\prime}(k_{2})=\beta(\pi,k_{2})\alpha(k_{2})\beta(-\pi,k_{2})^{-1}\quad\mbox{for }\beta\mbox{ such that }\quad J\beta(t,k_{2})=\overline{\beta(-t,-k_{2})}J.$
Then we can use the symmetry constraint on $\beta$ to rewrite
$\alpha^{\prime}(k_{2})=\beta(\pi,k_{2})\alpha(k_{2})J^{-1}\beta(\pi,-k_{2})^{t}J$
which shows how we can set
$\gamma_{0}^{\prime}:=\gamma_{0}J^{-1}\beta(\pi,0)^{t}J\quad\text{and}\quad\gamma_{\pi}^{\prime}:=\gamma_{\pi}J^{-1}\beta(\pi,\pi)^{t}J$
and obtain $\alpha^{\prime}(0)=J(\gamma_{0}^{\prime})^{t}J^{-1}\gamma_{0}^{\prime}$ and $\alpha^{\prime}(\pi)=J(\gamma_{\pi}^{\prime})^{t}J^{-1}\gamma_{\pi}^{\prime}$. Now, given a continuous function $b:[-\pi,\pi]^{2}\to\mathbb{R}$ such that $\det\beta(t_{1},t_{2})=e^{ib(t_{1},t_{2})}$, we have that $\lambda_{0}^{\prime}=\lambda_{0}+b(0,0)$, $\lambda_{\pi}^{\prime}=\lambda_{\pi}+b(\pi,0)$ and $\mu^{\prime}(k_{2})=\mu(k_{2})+b(\pi,k_{2})+b(\pi,-k_{2})$. Since $2\lambda_{0}^{\prime}=2\lambda_{0}+2b(\pi,0)=\mu^{\prime}(0)$, we can compute the invariant by studying
$2\lambda_{\pi}^{\prime}-\mu^{\prime}(\pi)=2\lambda_{\pi}+2b(\pi,\pi)-\mu(\pi)-b(\pi,\pi)-b(\pi,-\pi)=2\lambda_{\pi}-\mu(\pi)+b(\pi,\pi)-b(\pi,-\pi).$
According to Definition A.2, $b(\pi,\pi)-b(\pi,-\pi)]/2\pi$ is the winding number of $\det\beta(\pi,\cdot)]$ in the second direction. In view of Remark 4.2, this is even, so $b(\pi,\pi)-b(\pi,-\pi)\in 4\pi\mathbb{Z}$. We conclude that that
$e^{i\frac{2\lambda_{\pi}^{\prime}-\mu^{\prime}(\pi)}{2}}=e^{i\frac{2\lambda_{\pi}-\mu(\pi)}{2}}e^{i\frac{b(\pi,\pi)-b(\pi,-\pi)}{2}}=e^{i\frac{2\lambda_{\pi}-\mu(\pi)}{2}}.$
Therefore, the invariant is also well-posed with respect to the equivalence relation (9), as claimed. ∎
Remark 4.5.
The $\mathbb{Z}_{2}$ invariant introduced above can be viewed as a reformulation of the Graf–Porta invariant, first introduced in [graf2013bulk]. This reformulation is analogous to the one contained in Proposition 5.3 in [Cornean_2017], but extended to the case in which the projection-valued map $P$ has regularity $C^{0}$ rather than $C^{1}$. The above-mentioned Proposition shows that the Graf–Porta invariant $\mathcal{I}(\alpha)$ of a matching matrix $\alpha$ can be computed as the sum of twice the difference between the Pfaffians of $J\alpha(\pi)$ and $J\alpha(0)$ (which are anti-symmetric matrices), and the “partial winding number” of the determinant of the matching matrix $\alpha$ restricted to $0,\pi]$. This winding number is “partial” because on $0,\pi]$ the map $\alpha$ does not define a closed loop; however, in the $C^{1}$ setting it can still be computed as the integral, restricted to $0,\pi]$, of the derivative of $\log\det\alpha(k_{2})$.
In the $C^{0}$ setting, this derivative is not available, but the partial winding number can nonetheless be defined as the difference between the initial and final values of any continuous lift to the universal cover $\mathbb{R}$ of $\mathrm{U}(1)$ of the map $\det\alpha$ restricted to $0,\pi]$. Such a lift is simply a continuous logarithm of $\det\alpha(k_{2})$, since the covering map from $\mathbb{R}$ to $\mathrm{U}(1)$ is the exponential. In this context, the map $\mu(k_{2})$ defined earlier provides exactly such a logarithm, while the values $\lambda_{0}$ and $\lambda_{\pi}$ are the Pfaffians of $J\alpha(0)$ and $J\alpha(\pi)$, respectively. Since $\mu(0)=2\lambda_{0}$, we obtain
$2\lambda_{\pi}-\mu(\pi)=2(\lambda_{\pi}-\lambda_{0})\;-\;(\mu(\pi)-\mu(0))=2\operatorname{Pf}(J\alpha(\cdot))\big|_{0}^{\pi}\;+\;\log\det\alpha(\cdot)\big|_{0}^{\pi}.$
This, together with the discussion in [Cornean_2017], shows that the invariant defined in this work reproduces the Graf–Porta Index.
Using the invariant, the following theorem states that $A/\!\sim$ has only two elements, given by the equivalence classes identified by $\delta=1$ and $\delta=-1$. This result already appears in [Cornean_2017], but we include the proof —adapted to the present formalism— for completeness.
Theorem 4.6.
Given two families of symmetric matching matrices $\alpha,\alpha^{\prime}\in A$, the $\alpha\sim\alpha^{\prime}$ as in (9) if and only if $\delta(\alpha)=\delta(\alpha^{\prime})$.
Proof.
The previous result already showed the constancy of $\delta$ on $\sim$-equivalence classes, so we need to show that, conversely, if $\delta(\alpha)=\delta(\alpha^{\prime})\in\mathbb{Z}_{2}$ then it is possible to construct a function $\beta:[-\pi,\pi]\times S^{1}\to\mathrm{U}(2n)$ such that
(10) $\beta(t,k_{2})=J^{-1}\overline{\beta(-t,-k_{2})}J\mbox{ and }\alpha^{\prime}(k_{2})=\beta(\pi,k_{2})\alpha(k_{2})J^{-1}\beta(\pi,-k_{2})^{t}J.$
At $(t,k_{2})=(\pi,0)$ the equation on the right reduces to $\alpha^{\prime}(0)=\beta(\pi,0)\alpha(0)J^{-1}\beta(\pi,0)^{t}J$, which is satisfied for $\beta(\pi,0):=\gamma_{0}^{\prime}\gamma_{0}^{-1}$; the same holds at $(t,k_{2})=(\pi,-\pi)$ for $\beta(\pi,\pi):=\gamma_{\pi}^{\prime}\gamma_{\pi}^{-1}$. Since all $\gamma$’s can be chosen to be symplectic matrices, we have $\beta(\pi,0),\beta(\pi,\pi)\in\mathrm{Sp}(2n)$.
The next step is to extend the definition of $\beta$ to the segment $\overline{(\pi,-\pi)(\pi,0)}$. This can be achieved by choosing any continuous path $p\colon[0,1]\to\mathrm{U}(2n)$ with $p(0)=\beta(\pi,0)$ and $p(1)=\beta(\pi,\pi)$, which exists since $\mathrm{U}(2n)$ is path-connected, allowing us to define $\beta(\pi,k_{2})=p(-k_{2}/\pi)$ for $k_{2}\in[-\pi,0]$. To further extend the definition for $k_{2}\in[0,\pi]$, we impose the second condition in (10) and define
$\beta(\pi,s)=\alpha^{\prime}(s)J^{-1}\overline{\beta(\pi,-s)}J\alpha(s)^{-1}=\alpha^{\prime}(s)J^{-1}\overline{p(-s/\pi)}J\alpha(s)^{-1}\quad\forall s\in[0,\pi].$
Before moving on, we need to check that the winding number of $\det\beta(\pi,k_{2})$ is even, in order not to violate the condition in Remark 4.2. Using the same notation we used when we computed $\delta(\alpha^{\prime})$, we have that $\det\beta(\pi,0)=1=e^{i(\lambda_{0}^{\prime}-\lambda_{0})}$ and $\det\beta(\pi,\pi)=1=e^{i(\lambda_{\pi}^{\prime}-\lambda_{\pi})}$. Moreover, it is possible to chose a function $l:[0,1]\to\mathbb{R}$ such that $\det(p(s))=e^{il(s)}$ with $l(0)=\lambda_{0}^{\prime}-\lambda_{0},l(1)-\lambda_{\pi}^{\prime}+\lambda_{\pi}\in 2\pi\mathbb{Z}$ and use it to express
$\det\beta(\pi,k_{2})=\begin{cases}e^{il(-k_{2}/\pi)}&\mbox{if }k_{2}\in[-\pi,0],\\ e^{i[\mu^{\prime}(k_{2})-l(k_{2}/\pi)-\mu(k_{2})]}&\mbox{if }k_{2}\in[0,\pi],\end{cases}$
and finally $\det\beta(\pi,k_{2})=e^{ib(k_{2})}$ where
$b(k_{2})=\begin{cases}l(-k_{2}/\pi)&\mbox{if }k_{2}\in[-\pi,0],\\ \mu^{\prime}(k_{2})-l(k_{2}/\pi)-\mu(k_{2})&\mbox{if }k_{2}\in[0,\pi].\end{cases}$
The winding number of $\det(\beta(\pi,k_{2}))$ is now
$\det\beta(\pi,\cdot)]=\frac{\mu^{\prime}(\pi)-\mu(\pi)-l(1)-l(1)}{2\pi}.$
However $2l(1)\equiv-2\lambda_{\pi}^{\prime}+2\lambda_{\pi}\bmod 4\pi$, so
$\det(\beta(\pi,\cdot))]\equiv\frac{\mu^{\prime}(\pi)-2\lambda_{\pi}^{\prime}-\mu(\pi)+2\lambda_{\pi}}{2\pi}\bmod 2.$
But this quantity is even because
$e^{i\frac{\mu^{\prime}(\pi)-2\lambda_{\pi}^{\prime}-\mu(\pi)+2\lambda_{\pi}}{2}}=\delta(\alpha^{\prime})\delta(\alpha)=1.$
All these arguments ensure that the second condition of Equation (10) holds, so we move to the discussion of the first symmetry condition. At $t=0$, it reads $J\beta(0,k_{2})=\overline{\beta(0,-k_{2})}J$. If $\det(\beta(\pi,\cdot))]=2r$, then we can define
$\beta(0,k_{2}):=\operatorname{diag}(e^{irk_{2}},\underbrace{1,\cdots,1}_{n-1},e^{irk_{2}},\underbrace{1,\cdots,1}_{n-1})$
and it is trivial to check that $\beta(0,k_{2})$ satisfies the symmetry constraint and has $\det\beta(0,\cdot)]=2r=[\det\beta(\pi,\cdot)]$. By A.6 we can find a homotopy $\beta_{t}(k_{2})$ in $\mathrm{U}(2n)$ with $\beta_{0}(k_{2})=\beta(0,k_{2})$, $\beta_{1}(k_{2})=\beta(\pi,k_{2})$ and then define
$\beta(t,k_{2}):=\begin{cases}\beta_{t/\pi}(k_{2})&\mbox{for }t\in[0,\pi]\\ J^{-1}\overline{\beta_{-t/\pi}(-k_{2})}J&\mbox{for }t\in[-\pi,0]\end{cases}$
This map satisfies all the desired conditions and completes the proof of the Theorem. ∎
5. Results
In this Section, we present some features of a time-reversal symmetric projection-valued map $P$ on $\mathbb{T}^{2}$ which can be deduced from our definition 4.3 of the invariant $\delta(P)\in\mathbb{Z}_{2}$.
5.1. Symmetric splitting and pseudo-periodic symmetric frame
Through the next theorem, which constitutes the main result of this text, we relate the possible choices of pseudo-periodic symmetric frame for $P$ to choices of a symmetric splitting for it, in the sense of Definition 1.7.
Theorem 5.1.
Consider a time-reversal symmetric projection-valued map $P:\mathbb{T}^{2}\to\operatorname{Proj}_{2n}(\mathcal{H})$ and a pseudo-periodic symmetric frame $\{u_{j}(t,k_{2})\}_{j\in\{1,\cdots,2n\}}$ for it, with associated family of matching matrices $\alpha\colon\mathbb{T}^{1}\to\mathrm{U}(2n)$. For $h\in\mathbb{Z}$, the following statements are equivalent:
  1. (1)
    There exists a time-reversal symmetric splitting of $P(k)$ in the sense of Definition 1.7 with $\operatorname{Ch}(P^{-})=h$.
  2. (2)
    There exists a pseudo-periodic symmetric frame $\{v_{j}(t,k_{2})\}_{j\in\{1,\cdots,2n\}}$ in the sense of Definition 1.6 with pseudo-periodicity
    $v_{1}(\pi,k_{2})=e^{ihk_{2}}v_{1}(-\pi,k_{2}),\qquad v_{n+1}(\pi,k_{2})=e^{-ihk_{2}}v_{n+1}(-\pi,k_{2}).$
  3. (3)
    The matching matrices $\alpha$ are $\sim$-equivalent in AA to
    $\alpha^{\prime}(k_{2}):=\operatorname{diag}(e^{ihk_{2}},\underbrace{1,\cdots,1}_{n-1},e^{-ihk_{2}},\underbrace{1,\cdots,1}_{n-1})$
    (compare (9)).
  4. (4)
    $\delta(P)=(-1)^{h}\in\mathbb{Z}_{2}$.
Proof.
The first statement implies the second because, given a splitting $P(k)=P^{+}(k)+P^{-}(k)$, we can apply Theorem 3.5 to construct a pseudo-periodic frame $\{v_{j}(t,k_{2})\}_{j\in\{1,\cdots,n\}}$ of $P^{-}$, in the sense of Definition 3.1, with the first vector satisfying $v_{1}(\pi,k_{2})=e^{i\operatorname{Ch}(P^{-})k_{2}}v_{1}(-\pi,k_{2})$. Then we can then set $v_{n+j}(t,k_{2}):=Tv_{j}(-t,-k_{2})$ and define a pseudo-periodic frame for $P^{+}$. The union of the two collections will constitute a pseudo-periodic symmetric frame of $P$ in the sense of Definition 1.6.
Conversely, the second statement implies the first because, given such a pseudo-periodic frame, we can define
$P^{-}(t,k_{2}):=\sum_{1\leq j\leq n}\left|v_{j}(t,k_{2})\right\rangle\left\langle v_{j}(t,k_{2})\right|,\quad P^{+}(t,k_{2}):=\sum_{n+1\leq j\leq 2n}\left|v_{j}(t,k_{2})\right\rangle\left\langle v_{j}(t,k_{2})\right|.$
The symmetry condition on the frame imposes $TP^{+}(t,k_{2})=P^{-}(-t,-k_{2})T$ and it is obvious that $P^{+}(t,k_{2})P^{-}(t,k_{2})\equiv 0$. Moreover, the pseudo-periodic condition still implies
$\left|v_{j}(-\pi,k_{2})\right\rangle\left\langle v_{j}(-\pi,k_{2})\right|=\left|v_{j}(\pi,k_{2})\right\rangle\left\langle v_{j}(\pi,k_{2})\right|$
so we can interpret $P^{\pm}$ as functions of $k\in\mathbb{T}^{2}$. Moreover, the family of matching matrices associated to the frame $\{v_{j}(-\pi,k_{2})\}_{1\leq j\leq n}$ for $P^{-}$ is $\tilde{\alpha}(k_{2}):=\operatorname{diag}(e^{ihk_{2}},1,\ldots,1)\in\mathrm{U}(n)$, leading to $\operatorname{Ch}(P^{-})=h$ according to Definition 3.2.
The equivalence between the second and the third statements follows the discussion in Section 4.1.
Finally, the equivalence between the third and the fourth statements follows naturally from Theorem 4.6 and the computation of $\delta(\alpha^{\prime})$. In fact, it is immediate to check that $\det\alpha^{\prime}(k_{2})\equiv 1$, so we can take $\mu^{\prime}(k_{2})\equiv 0$; moreover, from $\alpha^{\prime}(0)=\mathds{1}$ we can chose $\gamma_{0}^{\prime}=\mathds{1}$ and $\lambda_{0}^{\prime}=0$. From
$\alpha^{\prime}(\pi)=\operatorname{diag}((-1)^{h},\underbrace{1,\cdots,1}_{n-1},(-1)^{h},\underbrace{1,\cdots,1}_{n-1}),$
it is possible to chose
$\gamma^{\prime}_{\pi}=\operatorname{diag}(i^{h},\underbrace{1,\cdots,1}_{n-1},i^{h},\underbrace{1,\cdots,1}_{n-1}).$
Since $\det(\gamma_{\pi}^{\prime})=(-1)^{h}$, we can choose $\lambda_{\pi}^{\prime}=h$. All these choices lead to $\delta(\alpha^{\prime})=(-1)^{h}$ and so we can conclude the proof of the Theorem. ∎
5.2. Addivity and stability of the $\mathbb{Z}_{2}$ invariant
The main consequence of Theorem 5.1 is that the $\mathbb{Z}_{2}$ invariant of a projection-valued map $P$ can be expressed through the Chern number of a projection-valued map $P^{-}$ for any symmetric splitting $P=P^{+}+P^{-}$. This relation between the two invariants is not new in the computational physics literature, and has been used to construct local markers for the Kane–Mele invariant [Gilardoni_2022, Ba__2024]. As already observed in [Ba__2024], the stability of such local markers depends on the spatial localization of the symmetric splitting, that is, in the present formalism, on the regularity of the associated projection-valued map. The previous result not only provides theoretical support for this numerical procedure, but also yields a method for constructing a sufficiently regular symmetric splitting.
Besides, Theorem 5.1 allows us to directly prove two fundamental properties of $\delta$: additivity and invariance with respect to the equivalence relations introduced in Definition 1.8. Both properties are already well known in the literature for other formulations of the $\mathbb{Z}_{2}$ invariant, as they arise naturally in geometric or $K$-theoretic formulations (e.g. [de2012topological, kellendonk2019cycliccohomologygradedcralgebras]), or can be proved explicitly in analytic formulations analogous to the one presented in this article [Cornean_2017].